OpenStudy (anonymous):

Help. I did a problem but my answer is not an option. Can you show me where I messed up?

OpenStudy (anonymous):

f[g(-4)]

OpenStudy (anonymous):

i ended up with f(3)=-4.

OpenStudy (anonymous):

and f(x) and g(x)???

OpenStudy (phi):

we need to know the definition of g(x) and f(x)

OpenStudy (phi):

the first step is replace x with -4 in the definition of g(x)

OpenStudy (anonymous):

gx i got 3 for. i just told you the answer of fx.

OpenStudy (phi):

do you have g(x)= expression with x in it ?

terenzreignz (terenzreignz):

No, I'm sure they meant what exactly are f(x) and g(x) ? Like, f(x) = 3x + 4 or something like that...

OpenStudy (anonymous):

f(x) = x2 − 13 and g(x) = x + 5

OpenStudy (phi):

g(x) = x + 5 g(-4) is short hand for "replace x with -4" g(-4)= -4 + 5 what do you get ?

OpenStudy (anonymous):

+1. then how did i get 3.

OpenStudy (phi):

now f(g(-4)) is f(1) f(x)= x^2-13 replace x with 1 what do you get ?

OpenStudy (anonymous):

-12. i still don't understand how i got -4 as my answer though. I did the work the same.

OpenStudy (phi):

It is hard to guess the musings of your mind. The important thing is do you see how to do this problem? It takes two steps: find g(-4) put that number in f(x) to find f(g(-4))

OpenStudy (anonymous):

Yeah. Thanks.